OpenStudy (anonymous):

Ginny's class ordered 3 pizzas. They ate five-twelfths of each pizza. How many pizzas did they eat?

OpenStudy (bahrom7893):

ANGRY BIRDS!! Nice prof pic! Very original, I beat the game with 3 stars on every level on my psp.. didn't get to collect all the golden eggs yet though.. I think..

OpenStudy (bahrom7893):

Now as for the problem

OpenStudy (bahrom7893):

Suppose class ordered 3 pizzas of the same size: 1 and 1 and 1 (let's call a pizza a 1) five twelfth of a pizza is: (5/12)*1 = 5/12. Now if they had 5/12 th of each pizzas then it would've been: (5/12)+(5/12)+(5/12) = 3*(5/12)

OpenStudy (bahrom7893):

That's the same as: 15/12 = 5/4.

OpenStudy (anonymous):

And So The Answer Is...

OpenStudy (bahrom7893):

5/4 is (4+1)/4 = (4/4)+(1/4) = 1 + (1/4), therefore they ate a pizza and a quarter of a pizza.